OpenStudy (anonymous):

Algebra 1 question (AWARDING MEDAL!)

OpenStudy (legends):

I can help:)

OpenStudy (anonymous):

solve each equation for the indicated variable. y = mx + b. solve for b

OpenStudy (anonymous):

Subtract nd from both sides, to get y-mx=b

OpenStudy (anonymous):

Oops! I meant mx.

OpenStudy (holly00d1248):

subtract mx from both sides to get y-mx=b

OpenStudy (anonymous):

can you write it out for me?

OpenStudy (holly00d1248):

like that